OpenStudy (anonymous):

what is the solution of x+y=7 2x-4y=8

jimthompson5910 (jim_thompson5910):

we have a `y` in the first equation and a `-4y` in the second equation adding them up gives: `y+(-4y) = -3y` which isn't 0 like we want however we can change that `y` to `+4y` by multiplying by 4. Then notice how `4y + (-4y) = 0y = 0`

OpenStudy (anonymous):

so it's 0?

jimthompson5910 (jim_thompson5910):

to change the y to 4y, we have to multiply both sides of the first equation by 4 x+y=7 4*(x+y)=4*7 4x+4y = 28

jimthompson5910 (jim_thompson5910):

This system x+y=7 2x-4y=8 turns into 4x+4y=28 2x-4y=8

jimthompson5910 (jim_thompson5910):

Do you see how to solve 4x+4y=28 2x-4y=8 ??

OpenStudy (anonymous):

i don't

OpenStudy (anonymous):

would you subtract?

jimthompson5910 (jim_thompson5910):

you add actually

jimthompson5910 (jim_thompson5910):

add straight down 4x+4y=28 2x-4y=8 ----------- 6x+0y = 36 so 6x = 36 which means x = ???

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

now take that x value and plug it into any equation that has x and y in it

jimthompson5910 (jim_thompson5910):

then solve for y

OpenStudy (anonymous):

12-47=8? am i doing that right?

OpenStudy (anonymous):

i mean 12-4y=8

jimthompson5910 (jim_thompson5910):

yes now solve for y

OpenStudy (anonymous):

is y 0?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

try again